OpenStudy (anonymous):

Help ASAp! Will fan and medal!! :) Polygons PQRST and VWXYZ are similar. If PQ = 6 feet, QR = 11 feet, RS = 2 feet, ST = 9 feet, TP = 9 feet, and XY = 4 feet, then what is the length of WX? 18 feet 6 feet 22 feet 40 feet

OpenStudy (imstuck):

Use similar polygon ratios here. Compare the sides given in each polygon. Segment XY in one is similar to RS in the other. WX in one is similar to QR in the other. Use those values to set up a ratio:\[\frac{ RS }{ XY }=\frac{ QR }{ WX }\]You have some lengths given to you so let's fill in for those:\[\frac{ 2 }{ 4}=\frac{ 11 }{ x }\]x is the length of WX, which is what they are asking you to find. Cross multiply and what do you get?

OpenStudy (imstuck):

Can you solve that for x? When you find it, that is the length of the missing side they are asking for.

OpenStudy (anonymous):

D ? @IMStuck

OpenStudy (imstuck):

2x = 44...x = 22. So it's C.